# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v3

@@ 0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
7 Scientific Frontiers for Accelerationist Ecologism
The accelerationist ecological thesis (Section˜6) holds that sustainability is best achieved through technological abundance rather than degrowth. In this section we provide rigorous scientific evidence across four frontiers, namely econometric decoupling, clean energy, matter creation, and biological engineering, demonstrating that the technologies for post-scarcity without ecological destruction are not speculative but empirically grounded.
7.1 Econometric Evidence: Technology Drives Environmental Improvement
The Environmental Kuznets Curve (EKC) hypothesis, first established by Grossman and Krueger (1995) and named by Panayotou (1993), proposes an inverted-U relationship between per capita income and environmental degradation: pollution rises during early industrialization but declines beyond a turning point as societies invest in cleaner technologies and environmental regulation. While Stern (2004) noted important statistical caveats, the most comprehensive systematic review to date, screening 11,500+ papers and analyzing 835 full texts, confirms that relative decoupling of GDP from resource use and emissions is frequent, and absolute decoupling has been achieved in several developed economies for both production-based and consumption-based CO2 emissions (Haberl et al., 2020).
The evidence for decoupling has strengthened dramatically. Le Quéré et al. (2019) analyzed 18 developed economies (representing 28% of global emissions) that achieved sustained emissions reductions between 2005 and 2015, finding that renewable energy deployment is the primary driver. The OECD Green Growth framework (OECD, 2011) documents that green productivity rose faster than traditional Total Factor Productivity across 30 member states. Way et al. (2022) demonstrate that technology cost forecasting based on Wright’s learning curves (Wright, 1936) and experience curves (Lafond et al., 2018) accurately predicts the cost decline trajectory of clean energy technologies, and that a fast transition to renewables would save the global economy approximately \$12 trillion compared to maintaining fossil fuel systems.
Critically, Acemoglu et al. (2012) provide the theoretical foundation for directed technical change: policy interventions that redirect innovation toward clean technologies can permanently alter the trajectory of economic growth away from pollution-intensive paths. Combined with McAfee (2019)’s demonstration of “dematerialization,” the empirical finding that advanced economies are using fewer total resources even as GDP grows, and the World Bank’s framework for inclusive green growth (World Bank, 2012), the econometric evidence strongly supports the accelerationist thesis: the path to ecological sustainability runs through technological acceleration, not degrowth.
7.2 Frontier Clean Energy: Toward Unlimited Clean Power
The convergence of multiple breakthrough energy technologies makes the prospect of abundant, clean, decentralized energy credible for the first time in history.
In nuclear fusion, critical milestones have been crossed. The National Ignition Facility (NIF) achieved target gain $Q>1$ in December 2022, producing 3.15 MJ of fusion energy from 2.05 MJ of laser input, the first laboratory demonstration of fusion ignition (Abu-Shawareb et al., 2024). The SPARC tokamak at MIT, using high-temperature superconducting (REBCO) magnets that achieved a record 20-tesla field, projects $Q\approx 11$ with operations beginning in 2026 (Creely et al., 2020). The Wendelstein 7-X stellarator demonstrated reduced neoclassical energy transport, validating the optimized stellarator concept with record triple product values (Beidler et al., 2021), and Wurzel and Hsu (2022) provide a unified framework for assessing fusion progress across all approaches against the Lawson criterion.
In solar photovoltaics, perovskite-silicon tandem cells have shattered efficiency records. LONGi achieved NREL-certified 33.89% power conversion efficiency in 2024, rising to 34.85% in 2025, the first tandem cell to exceed the single-junction Shockley-Queisser limit (Liu et al., 2024; Green et al., 2024). These cells can be printed on flexible substrates, enabling solar integration into virtually any surface at dramatically reduced manufacturing costs.
Enhanced Geothermal Systems (EGS) have moved from concept to commercial deployment. Fervo Energy demonstrated sustained power generation from engineered reservoirs, with a 400 MW commercial plant under construction (Tester et al., 2006). The US DOE estimates 5,500+ GW of EGS potential in the United States alone, enough to power the entire US economy multiple times over, providing 24/7 carbon-free baseload energy accessible virtually anywhere on Earth.
In multi-day energy storage, iron-air batteries have been revolutionary. Form Energy’s 100+ hour storage technology at projected costs below \$20/kWh uses only iron, water, and air, some of Earth’s most abundant and non-toxic materials (Kwon et al., 2024). Combined with green hydrogen produced at 95%+ efficiency via innovations like H2Pro’s E-TAC electrolysis (Chatterjee et al., 2021), and DOE-backed long-duration storage programs (U.S. Department of Energy, 2024), the intermittency problem of renewables is being solved.
7.3 Creating Matter from Energy: Post-Scarcity Manufacturing
The relationship $E=mc^{2}$ implies that matter and energy are interconvertible. While wholesale matter creation from pure energy remains energetically impractical at current scales, the frontier of advanced manufacturing is converging toward the ability to create virtually any material object from basic feedstocks with minimal waste.
Pair production, the creation of matter-antimatter pairs from photon collisions, has been experimentally demonstrated. Breit and Wheeler (1934) provided the theoretical prediction, Burke et al. (1997) achieved the first laboratory demonstration at SLAC, and the STAR Collaboration at RHIC provided definitive evidence of the linear Breit-Wheeler process with real photons (STAR Collaboration, 2021). While pair production remains a physics demonstration rather than a manufacturing method, it establishes the fundamental principle that matter can be created from energy.
Additive manufacturing is achieving practical matter creation from digital designs. Metal 3D printing achieves buy-to-fly ratios of 1:1 to 3:1, compared to 10:1–30:1 for conventional machining, dramatically reducing material waste (DebRoy et al., 2018). 3D bioprinting fabricates living tissues including skin, bone, vascular grafts, and cartilage (Murphy and Atala, 2014). DNA origami enables atomically precise construction at the nanoscale (Rothemund, 2006), progressing toward Drexler’s vision of molecular manufacturing (Drexler, 1992).
Nuclear transmutation has created entirely new elements that do not exist in nature. Using 48Ca beams on actinide targets, researchers at JINR, GSI, and RIKEN have synthesized elements 107–118 (Oganessian and Utyonkov, 2015), demonstrating that humans can convert one form of matter into another using kinetic energy, genuine transmutation at industrial research scales.
7.4 Biological Engineering: Super Plants and Living Factories
The engineering of photosynthesis represents one of the most promising frontiers for simultaneously addressing energy, carbon capture, and food production.
The RIPE Project (Realizing Increased Photosynthetic Efficiency) has achieved remarkable results. South et al. (2019) engineered a synthetic photorespiratory bypass in tobacco that increased field biomass productivity by over 40%, the largest photosynthetic improvement ever demonstrated in field trials. Kromdijk et al. (2016) accelerated recovery from photoprotection, increasing biomass by 14–20% through a mechanism common to all flowering plants. The C4 Rice Project at IRRI (Ermakova et al., 2020) aims to install C4 photosynthetic machinery into rice, potentially increasing efficiency by 50% in tropical conditions. Salesse-Smith et al. (2018) achieved a 30% increase in RuBisCO content and 15% increase in CO2 assimilation through enzyme overexpression in maize.
[…]
A particularly striking frontier is biophotovoltaics (BPV), which harness photosynthetic organisms to generate electricity directly. McCormick et al. (2015) review the field, documenting how cyanobacteria and algae transfer electrons extracellularly to electrode surfaces. Plant-microbial fuel cells (Strik et al., 2008) generate electricity from living plant root exudates, with the Plant-e technology already commercialized. Engineering CAM photosynthesis into C3 crops (Borland et al., 2014) could reduce water use by up to 80%, enabling food and bioenergy production on arid lands.
7.5 Complementary Sciences: From Quantum Computing to Circular Economy
Several additional scientific frontiers reinforce the accelerationist ecological thesis.
In quantum computing for materials design, the variational quantum eigensolver (Peruzzo et al., 2014) and quantum chemistry algorithms (Cao et al., 2019) enable exponentially faster simulation of molecular behavior, potentially revolutionizing the rational design of catalysts, battery materials, and superconductors without costly trial-and-error synthesis. In nanotechnology, from carbon nanotubes (Iijima, 1991) to graphene (Novoselov et al., 2004), nanomaterials offer revolutionary capabilities for energy storage, water purification, and structural materials at a fraction of the weight and environmental cost of conventional alternatives. Nature’s 3.8 billion years of evolution provide what biomimicry researchers call open-source design blueprints for efficient buildings, drag-reducing surfaces, and ultra-tough materials (Benyus, 1997), democratizing innovation without intellectual property barriers.

### v5

@@ 0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
The apparent opposition between degrowth and accelerationism is a false dichotomy when viewed from the perspective of anarchist pacifism. Both movements share the same foundational commitment: opposition to capitalist domination, commitment to ecological sustainability, and pursuit of human flourishing through non-hierarchical social organization. The difference lies in technological strategy, not political values. This can be formalized as a multi-objective optimization problem. Define two objectives: $f_{1}$, the ecological risk during transition (probability of catastrophic environmental damage from betting on immature technologies), and $f_{2}$, the human suffering during transition (poverty, deprivation, and inequality persisting until post-scarcity technologies mature). Digital ecological degrowth occupies one region of the Pareto front: lower ecological risk (fewer technological gambles, reduced material throughput) but higher and more prolonged human suffering (because post-scarcity technologies arrive later, if at all, and material deprivation continues in the interim). Peaceful anarcho-accelerationism occupies the other region: higher ecological risk (betting on frontier technologies whose environmental impacts are not fully characterized) but lower human suffering (because technological abundance arrives sooner, relieving material deprivation more rapidly). No point on this front dominates the other; both positions are Pareto-optimal given different weightings of the two objectives. Figure˜8 illustrates this relationship.
DegrowthAccelerationismAnarchist pacifism(shared substrate)Tech. progressCurrent Pareto frontShifted frontEcological risk during transitionHuman suffering during transition
Figure 8: The degrowth-accelerationism tradeoff as a Pareto front. Both positions share anarchist pacifism as substrate and are Pareto-optimal given different weightings of ecological risk vs. human suffering. Technological progress (dashed curve) shifts the entire front toward the origin, reducing both risk and suffering. A mature movement supports both strategies simultaneously.
The critical insight is that the Pareto front is not fixed. Technological progress shifts it. Every advance in clean energy, every improvement in recycling efficiency, every breakthrough in sustainable manufacturing moves the entire front toward the origin: less risk and less suffering simultaneously. Accelerationism, by investing in these technologies, shrinks the Pareto front itself. This reframing transforms the debate: degrowthists and accelerationists are allies sharing anarchist pacifist foundations (Kallis, 2018; Hickel, 2020; Latouche, 2009), differing only in their position along a Pareto-optimal tradeoff. A mature movement would support both strategies simultaneously, hedging ecological bets through conservation and sufficiency (degrowth wisdom) while accelerating clean technology deployment (accelerationist commitment). What is fundamental and shared is the anarchist pacifism that underlies both positions: the rejection of domination, the commitment to commons governance, and the insistence on peaceful means. The technological strategy is secondary to the political ethics.
The Liberation Stack is not merely a list of technologies; it is a dependency graph with explicit preconditions for each layer. No layer can be sustainably deployed without the layers below it in the dependency ordering. This section makes the preconditions and gate conditions explicit, so that communities attempting to build the stack can evaluate their readiness for each layer and plan their implementation sequence rationally. Table˜9 summarizes the dependency structure.
Table 9: Liberation Stack preconditions and gate conditions. Each layer requires specific preconditions to be met before sustainable deployment. The dependency column indicates which lower layers must be operational.
Layer Domain Dependencies Gate condition for deployment
0 Energy None Community achieves >>50% renewable self-generation capacity
1 Manufacturing Energy Access to clean energy + open-source hardware designs for local fabrication
2 Food Energy, partial Manufacturing Energy for controlled environment agriculture + fabrication of farming tools and sensors
3 Communication Energy Electricity for servers, routers, and mesh nodes; can deploy in parallel with Manufacturing
4 Knowledge Communication Connectivity for accessing and sharing knowledge commons; functional federated platforms
5 Governance Communication, Knowledge Informed, connected citizens capable of sustained deliberation; digital literacy
Energy is the precondition for everything. Without clean energy sovereignty, no other layer achieves long-term sustainability: manufacturing depends on power for CNC machines and 3D printers, food production depends on power for irrigation pumps and controlled-environment systems, communication depends on electricity for servers and routers, and governance depends on the infrastructure that communication provides. The gate condition for the energy layer is that a community must achieve greater than 50% renewable self-generation capacity before the energy layer can be considered operational. This threshold is not arbitrary: below 50%, the community remains structurally dependent on external energy providers and vulnerable to supply disruptions or price manipulation. Current examples demonstrate that this gate is achievable. The World Bank is accelerating renewable deployment in developing countries through its Climate Change Action Plan, with particular emphasis on distributed solar in Sub-Saharan Africa and South Asia. The European Union’s Clean Energy Package includes directives enabling energy communities to generate, store, and sell renewable energy collectively, with over 3,400 energy cooperatives now operational across Europe (Kunze and Becker, 2015). Community solar programs in the United States have grown to over 6 GW of capacity, providing clean energy access to households that cannot install rooftop panels. The Just Energy Transition Partnerships (JETPs) negotiated by the World Bank and partner governments in South Africa, Indonesia, and Vietnam (World Bank, 2023) represent the most ambitious international framework yet for financing the transition from fossil fuels to renewables in developing economies, with combined pledges exceeding \$40 billion. A significant consideration within the energy layer is nuclear waste management. Conventional approaches to nuclear waste storage, including deep geological repositories, face both technical challenges and public opposition. An unconventional but increasingly plausible solution is space disposal via reusable rockets: as launch costs decline from approximately \$10,000/kg (Space Shuttle era) to projections below \$100/kg with fully reusable vehicles such as SpaceX Starship, the economics of launching vitrified high-level waste into solar orbit or toward the Sun become tractable. We note this not as a near-term proposal but as an illustration of how technological acceleration opens solution spaces that were previously dismissed as economically absurd.
Manufacturing requires the energy layer as a precondition. The gate condition is access to clean energy combined with open-source hardware designs sufficient for local fabrication of essential goods. Without energy, fablabs are expensive hobbies rather than productive infrastructure. With energy, the global network of 2,500+ fablabs (Fab Foundation, 2023) demonstrates that distributed manufacturing is operational and can produce everything from medical devices to housing components. Food sovereignty requires energy and partial manufacturing capability. The gate condition is energy for controlled-environment agriculture (vertical farms, greenhouses, aquaponics) combined with the capacity to fabricate farming tools, sensors, and irrigation systems locally. Communication requires only energy as a prerequisite and can be deployed in parallel with manufacturing. The gate condition is electricity sufficient to power servers, routers, and mesh nodes. The guifi.net model with 40,000+ nodes (Baig et al., 2015) demonstrates that community communication infrastructure can be built incrementally, with each new node extending the network. The Fediverse demonstrates that application-layer communication can be federated without centralized control (Zulli et al., 2020). Knowledge requires communication as a prerequisite: without connectivity, knowledge commons cannot be accessed or shared. The gate condition is functional federated communication platforms with sufficient bandwidth for knowledge exchange. Governance requires both communication and knowledge: informed deliberation presupposes access to information and the ability to communicate. The gate condition is a connected, informed citizenry with sufficient digital literacy to participate in online deliberative processes through tools such as Decidim (Barandiaran et al., 2024), Pol.is (Small et al., 2021), and Loomio.
The preconditions framework has a critical implication for the relationship between the Liberation Stack and the state. The framework proposes not revolution but progressive state dissolution: a gradual, layer-by-layer process in which the state’s functions are incrementally assumed by commons-based governance as each layer demonstrates competence. This process is explicitly incremental, reversible at every stage, and compatible with existing democratic institutions. It is not a seizure of state power; it is the progressive construction of alternatives that make state functions redundant.
In the first phase, the state remains fully operational but creates legal space for commons institutions. Energy cooperatives operate under existing cooperative law. Platform cooperatives register as conventional businesses. Community mesh networks operate under telecommunications regulations. The state’s role is to provide the legal framework within which commons institutions can form, experiment, and demonstrate viability. This phase requires no changes to constitutional order and is already underway in many jurisdictions: the EU energy communities directive, cooperative promotion legislation in most democracies, and digital commons policies in cities like Barcelona provide the legal infrastructure.
In the second phase, commons institutions provide an increasing share of public services, and the state’s role shifts from direct provider to guarantor of rights. Community energy cooperatives provide a growing fraction of the energy supply. Platform cooperatives handle increasing shares of local economic activity. Federated communication platforms provide alternatives to corporate social media. The state retains its regulatory and redistributive functions but delegates operational service delivery to commons institutions that have demonstrated competence. The key assumption is that commons institutions must demonstrate superior performance on measurable outcomes (cost, quality, accessibility, ecological impact) before the state delegates further functions to them.
In the third phase, federated commons governance handles the majority of resource allocation within participating communities, and the state’s role reduces to conflict resolution between federations and external defense. The governance layer of the Liberation Stack provides the institutional infrastructure for democratic self-governance at multiple scales: local communes for neighborhood decisions, regional federations for resource sharing, and inter-regional coordination for issues that transcend local boundaries. The state’s residual functions are those that require coercive authority: enforcement of ecological boundaries, protection of individual rights against community tyranny, and defense against external aggression.
In the fourth phase, the state withers as commons governance matures, and international federation handles coordination between self-governing regions. This phase is the most speculative and the most distant. It requires not only the successful operation of all Liberation Stack layers at regional scale but also the development of effective inter-regional governance protocols, the resolution of free-rider problems between federations, and the management of common-pool resources (atmosphere, oceans, biodiversity) at planetary scale. We present this phase as a theoretical horizon, not as a near-term prediction. The critical precondition for each phase is demonstrated competence at the previous level: if commons governance fails at local scale, there is no expansion to regional scale. If regional federations cannot coordinate effectively, there is no progression to inter-regional governance. This constraint is not merely prudential; it is structural. A commons governance system that has not proven itself at small scale has no basis for claiming authority at larger scale. The reversibility of each phase is essential: if a commons institution fails to deliver on its promises, the state’s functions can be restored without systemic collapse. This is the fundamental difference between progressive state dissolution and revolutionary state seizure: the former is incremental and reversible, the latter is sudden and irreversible.
The accelerationist ecological thesis is supported not only by the energy technologies already discussed but by a broader set of scientific frontiers that make the prospect of material abundance within ecological constraints increasingly credible. We survey these frontiers across four domains: econometric evidence for decoupling, biological engineering, advanced manufacturing toward post-scarcity, and complementary sciences.
The econometric evidence for the compatibility of technological progress with environmental sustainability has strengthened dramatically. The Environmental Kuznets Curve (EKC) hypothesis (Grossman and Krueger, 1995; Panayotou, 1993) proposes an inverted-U relationship between per capita income and environmental degradation: pollution rises during early industrialization but declines beyond a turning point as societies invest in cleaner technologies and environmental regulation. While Stern (2004) noted important statistical caveats regarding the generalizability of the EKC across pollutants and countries, the most comprehensive systematic review to date, screening 11,500+ papers and analyzing 835 full texts, confirms that relative decoupling of GDP from resource use and emissions is frequent, and absolute decoupling has been achieved in several developed economies for both production-based and consumption-based CO2 emissions (Haberl et al., 2020). The evidence has continued to strengthen. Le Quéré et al. (2019) analyzed 18 developed economies representing 28% of global emissions that achieved sustained emissions reductions between 2005 and 2015, finding that renewable energy deployment is the primary driver of decarbonization. The OECD Green Growth framework (OECD, 2011) documents that green productivity rose faster than traditional Total Factor Productivity across 30 member states. Way et al. (2022) demonstrate that technology cost forecasting based on Wright’s learning curves (Wright, 1936) and experience curves (Lafond et al., 2018) accurately predicts the cost decline trajectory of clean energy technologies, and that a fast transition to renewables would save the global economy approximately \$12 trillion compared to maintaining fossil fuel systems. Critically, Acemoglu et al. (2012) provide the theoretical foundation for directed technical change: policy interventions that redirect innovation toward clean technologies can permanently alter the trajectory of economic growth away from pollution-intensive paths. Combined with McAfee (2019)’s demonstration of dematerialization, the empirical finding that advanced economies are using fewer total resources even as GDP grows, and the World Bank’s framework for inclusive green growth (World Bank, 2012), the econometric evidence strongly supports the accelerationist thesis: the path to ecological sustainability runs through technological acceleration, not degrowth.
Biological engineering represents one of the most promising frontiers for simultaneously addressing energy, carbon capture, and food production. The RIPE Project (Realizing Increased Photosynthetic Efficiency) has achieved remarkable results: South et al. (2019) engineered a synthetic photorespiratory bypass in tobacco that increased field biomass productivity by over 40%, the largest photosynthetic improvement ever demonstrated in field trials. Kromdijk et al. (2016) accelerated recovery from photoprotection, increasing biomass by 14–20% through a mechanism common to all flowering plants, suggesting broad applicability across crop species. The C4 Rice Project at IRRI (Ermakova et al., 2020) aims to install C4 photosynthetic machinery into rice, potentially increasing efficiency by 50% in tropical conditions where rice is a staple for billions. Salesse-Smith et al. (2018) achieved a 30% increase in RuBisCO content and 15% increase in CO2 assimilation through enzyme overexpression in maize. Synthetic biology has created artificial photosynthetic systems that surpass natural ones: Miller et al. (2020) constructed an artificial chloroplast by encapsulating spinach photosynthetic membranes in cell-sized droplets powering the synthetic CETCH cycle, achieving CO2 fixation rates 100 times faster than previous approaches. The CETCH cycle itself (Schwander et al., 2016), comprising 17 enzymes from 9 organisms, uses an enzyme (CCR) that fixes CO2 approximately 20 times faster than plant RuBisCO. A particularly striking frontier is biophotovoltaics (BPV), which harness photosynthetic organisms to generate electricity directly. McCormick et al. (2015) review the field, documenting how cyanobacteria and algae transfer electrons extracellularly to electrode surfaces, producing electricity from sunlight and water with no fuel input. Plant-microbial fuel cells (Strik et al., 2008) generate electricity from living plant root exudates, with the Plant-e technology already commercialized for small-scale applications. Engineering CAM (Crassulacean Acid Metabolism) photosynthesis into C3 crops (Borland et al., 2014) could reduce water use by up to 80%, enabling food and bioenergy production on arid lands that are currently considered unsuitable for agriculture.
The frontier of advanced manufacturing is converging toward the ability to create virtually any material object from basic feedstocks with minimal waste. The relationship $E=mc^{2}$ implies that matter and energy are interconvertible; while wholesale matter creation from pure energy remains energetically impractical at current scales, pair production, the creation of matter-antimatter pairs from photon collisions, has been experimentally demonstrated. Breit and Wheeler (1934) provided the theoretical prediction, Burke et al. (1997) achieved the first laboratory demonstration at SLAC, and the STAR Collaboration at RHIC provided definitive evidence of the linear Breit-Wheeler process with real photons (STAR Collaboration, 2021). While pair production remains a physics demonstration rather than a manufacturing method, it establishes the fundamental principle that matter can be created from energy. In practical manufacturing, additive manufacturing is achieving near-zero-waste production: metal 3D printing achieves buy-to-fly ratios of 1:1 to 3:1, compared to 10:1 to 30:1 for conventional machining, dramatically reducing material waste (DebRoy et al., 2018). 3D bioprinting fabricates living tissues including skin, bone, vascular grafts, and cartilage (Murphy and Atala, 2014). DNA origami enables atomically precise construction at the nanoscale (Rothemund, 2006), progressing toward Drexler’s vision of molecular manufacturing (Drexler, 1992). Nuclear transmutation has created entirely new elements that do not exist in nature: using 48Ca beams on actinide targets, researchers at JINR, GSI, and RIKEN have synthesized elements 107–118 (Oganessian and Utyonkov, 2015), demonstrating that humans can convert one form of matter into another using kinetic energy at industrial research scales.